OpenStudy (anonymous):

What is the equation of the line that passes through (2, 1) and is perpendicular to the graph of y = –x + 2? y = x – 1 y = x + 1 y = –x + 3 y = –x + 2

jimthompson5910 (jim_thompson5910):

The slope of y = -x+2 is -1

jimthompson5910 (jim_thompson5910):

so the slope of any perpendicular line is 1 (flip the fraction, flip the sign)

jimthompson5910 (jim_thompson5910):

y = mx + b y = 1*x + b ... plug in the perpendicular slope 1 = 1*2 + b ... plug in the given point 1 = 2 + b 1 - 2 = b -1 = b b = -1 So the equation of the perpendicular line is y = x - 1

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

np